When we want confidence and use the conservative estimate of 95%, we can use the simple formula to roughly determine the sample size needed for a given margin of error p=.05. Use this formula to determine the sample size needed for a margin of error of 0.01 .
Answer #1

Solution :

Given that,

\hat p = 0.05

1 - \hat p = 0.95

margin of error = E = 0.01

Z\alpha/2 = Z0.025 = 1.96

sample size = n = (Z\alpha / 2 / E )2 * \hat p * (1 - \hat p )

= (1.96 / 0.01)2 * 0.05 * 0.95

= 1825

sample size = n = 1825